the potential difference on the surface of charged sphere radius R and between the distance 3r from the surface of a sphere is v the electric field at the distance which is equal to radius r from the surface will be

Grade:12th pass

1 Answers

Khimraj
5 years ago
potential at surface = KQ/r
potential at distance 3r from the surface = KQ/4r
KQ/r – KQ/4r = v
3KQ/4r = v
KQ/r = 4v/3
electric field at the distance which is equal to radius r from the surface = KQ/4r2 = (KQ/r)/(4r) = (4v/3)/(4r) = v/3r
Hope it clears............................